OpenStudy (anonymous):

please help Keith started saving for retirement at age 45 with plans to retire at age 70. He invested an average of $500 per month in various securities, with an average annual return of 6% adjusted for inflation. Assuming monthly compounding, how much has Keith saved at the start of retirement? $27,432.26 $446,496.98 $346,496.98 $712,355.15

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

can you explain!

OpenStudy (mathstudent55):

\( FV = A \times \dfrac {(1 + i)^n - 1}{i} \) FV = future value n = number of periods i = interest rate per period as a decimal A = annualized (monthly) payment

OpenStudy (mathstudent55):

\( FV = A \times \dfrac {(1 + i)^n - 1}{i}\) \(n = 25 \times 12\) = 300 \(i = \dfrac {6}{12 \times 100} = 0.005\) \( FV = $500 \times \dfrac {(1 + 0.005)^{300} - 1}{0.005} \) \(FV = $346,496.98 \) Choice C.